OpenStudy (anonymous):

what is the sum of the first 30 terms of an=3n+2

OpenStudy (anonymous):

what is the first term, if \(n=1\) ?

OpenStudy (anonymous):

there is no first term

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what do you get if you replace \(n\) by \(1\) in \(a_n=3n+2\) ?

OpenStudy (anonymous):

5 ?

OpenStudy (anonymous):

right, so the first term is \(a_1=5\) what is the last term, if \(n=30\) ?

OpenStudy (anonymous):

92 ?

OpenStudy (anonymous):

yes what do you get when you add them?

OpenStudy (anonymous):

i mean when you add \(a_1+a_{30}=5+92\) ?

OpenStudy (anonymous):

97

OpenStudy (anonymous):

yes last step is to take half of 30, which is 15,and multiply \(15\times 97\)

OpenStudy (anonymous):

so 1455 ?

OpenStudy (anonymous):

and would you be able to help with something else ?

OpenStudy (anonymous):

formula looks like \[S=\frac{n}{2}\left(a_1+a_n\right)\] which in your case is \[\frac{30}{2}(5+92)=15\times 97\]

OpenStudy (anonymous):

sure i have one minute only though

OpenStudy (anonymous):

if you save three pennies on january 1, six on january 2, 9 on january 3 and continue for 1 year not a leap year, what will be the value of your entire serving in dollars at the end of 1 year

OpenStudy (anonymous):

even that was wrong, let me try again

OpenStudy (anonymous):

\[3+2\times 3+3\times 3+...+364\times 3\] \[a_1=3, a_n=3n\] first term is \(3\) last term is \(365\times 3\) add up \[\frac{365}{2}(3+365\times 3)\]

OpenStudy (anonymous):

so 133772.5 ?

OpenStudy (anonymous):

now i got $2014.80